# Université de Genève, Genève, CH-1211 Switzerland

### v2

@@ 2.1 The A model
see Fig. 1 for a pictorial representation. Let $[S_{i}]\in H_{2}(M,{\mathbb{Z}})$, $i=1,\cdots,s$, be a basis for the two-homology of $M$, with $s=b_{2}(M)$ as before. The maps (2.3) are classified topologically by the homology class
@@ 2.2 The Gopakumar–Vafa representation
Note that, due to the vanishing property of the GV invariants mentioned above, the sum over $g$ in (2.19) is finite. One could think that (2.18) can perhaps be summed, as a series now in the variables ${\rm e}^{-t_{i}}$. It turns out that the properties of this series depend crucially on the value of $g_{s}$. If $g_{s}$ is real, the series is not even well-defined. This is due to the inverse square sines appearing in (2.14), which lead to singularities at rational values of $g_{s}$. In fact, given any rational value of $g_{s}$, there is a minimum degree ${\bf m}_{\rm min}$ such that infinitely many coefficients $F_{\bf m}(g_{s})$ with ${\bf m}>{\bf m}_{\rm min}$ are singular at that rational value. As a consequence, given any real value of $g_{s}$, rational or not, there is a degree starting from which infinitely many coefficients $F_{\bf m}(g_{s})$ can be made arbitrarily large. This is clearly a pathological situation. One could still try to make sense of the r.h.s. of (2.18) for complex values of $g_{s}$. However, in that case there is evidence that the coefficients $F_{\bf m}(g_{s})$ grow with Gargantuan speed, even worst than factorial. For example, in the case of local ${\mathbb{P}}^{2}$, introduced in (2.2), one has $b_{2}(M)=1$ (corresponding to the ${\mathbb{P}}^{1}$ inside ${\mathbb{P}}^{2}$) and therefore there is a single degree. Explicit computations suggest that hmmo ; mmrev
@@
There is a special class of toric CY manifolds which can be used to engineer five-dimensional $SU(N)$ gauge theories kkv ; n . For these manifolds, the gauge theory instanton partition function (reviewed in N. Nekrasov’s lectures in this school) provides a rearrangement of the GV expansion which converges for complex $g_{s}$ felder-lennert . However, it is still singular if $g_{s}\in 2\pi\mathbb{Q}$.
@@ 2.4 The B model
[…]
@@
The prepotential gives the genus zero free energy of the topological string, in the B model. An important bonus of using the B model is that the genus zero free energy is obtained as a global function on the moduli space of the CY manifold.
@@
$\widetilde{F}_{0}(\tilde{t})=F_{0}(t)-S(t,\tilde{t}),$ (2.39)
which is a generalized Legendre transform. The function $S(t,\tilde{t})$ has the form
@@
The coefficients appearing in this polynomial can be related to the parameters appearing in (2.37) by imposing that $\widetilde{F}_{0}(\tilde{t})$ is independent of $t$,
@@
where the function $S(t,\tilde{t})$ implementing the transform is given by (2.43). The integral in the r.h.s. of (2.44) has to be understood formally, since the total free energy appearing in the integrand is itself a formal power series. To obtain the transformation properties of the genus $g$ free energies, we evaluate the integral in the r.h.s. of (2.44) in a saddle point approximation for $g_{s}$ small. At leading order we recover the generalized Legendre transform (2.39), and in particular the condition for a saddle point is precisely (2.41). The evaluation at higher orders leads to explicit transformation properties for the higher genus free energies, see abk for examples.
@@ 2.5 A more complicated example: local ${\mathbb{P}}^{2}$
Figure 2: The moduli space of local ${\mathbb{P}}^{2}$, with the three special points $z=0$ (large radius), $z=-1/27$ (conifold) and $z=\infty$ (orbifold).
@@
It can be seen from the PF equations that the convergence radius of the series (2.54) at $z=0$ is set by the conifold singularity at $z=-1/27$. This means in particular that the radius of convergence of the genus zero free energy (2.57) is set by the value of $t$ at the conifold point cdgp ; kz , which in this case is given by villegas
@@
The second relation is of course a consequence of (2.59). The choices of signs in (2.61) are correlated with the choice of branch cut of $\log(z)$ for $-1/27<z<0$, and they have been made in such a way that $\lambda$ and $\partial_{\lambda}F_{0}^{c}$ are real in that interval. The genus zero free energy in the conifold frame can be computed explicitly and it has the form,
@@ 3.1 Warm-up: the conifold and the resolved conifold
$\0\zeta=2\pi{\rm i}\ell\lambda,\qquad\ell\in{\mathbb{Z}}\backslash\{0\},$ (3.3)
@@
$\0\varphi_{\ell{\cal A}}(g_{s})={1\over 2\pi}\left\{{1\over\ell}\left({{\cal A}\over g_{s}}\right)+{1\over\ell^{2}}\right\}{\rm e}^{-\ell{\cal A}/g_{s}},\qquad\ell\in{\mathbb{Z}}\backslash\{0\}.$ (3.6)
@@
Therefore, the singularities of the Borel transform are of the form $\ell{\cal A}_{m}$, where \$\0\ell\in{\mathbb{Z}}\backslash\{0\}$ and the “action” ${\cal A}_{m}$ is labelled by an additional integer $m$:
@@ 3.2 The general multi-instanton trans-series
$S(\tilde{t},t)=-{t\tilde{t}\over c}+{d\over 2c}t^{2}+{4\pi^{2}{\rm i}m\over c}t+s(\tilde{t}),$ (3.22)
@@
so (3.32) is manifestly a non-perturbative correction. The exponent in (3.35) can be interpreted as the difference between the free energies of two different backgrounds, or points in the moduli space of the CY: the background $t$, and the background $t-cg_{s}$, in which $t$ is shifted. It is easy to see that in the $\ell$-th instanton sector the shift is given by $t-\ell cg_{s}$. Since, with the appropriate normalizations, $c$ is an integer, this suggests that $t$ is “quantized” in units of $g_{s}$. Such a quantization is typical of topological string theories with large $N$ duals, in which the CY modulus is interpreted as a ’t Hooft parameter and has the form $Ng_{s}$, where $N$ is the rank of the matrix model dv ; gv-cs ; mz . We will elaborate on this in section 4.
@@ 3.3 BPS states and the resurgent structure of topological strings
[…]
@@
where summation over the repeated indices is understood. Let us note that, in the case of toric CY manifolds, D6 branes decouple, and the charge ${\boldsymbol{\gamma}}$ is specified by $12s+$ 1 integers which we will denote by $c^{a}$, $d_{a}$ and $m$, with $a=1,\cdots,s$. The central charge reads then
@@
See Fig. 5 for a plot of the curve of marginal stability in the $u$-plane. States with charges $-{\boldsymbol{\gamma}}$ also belong to the spectrum.
@@
    The singularities display a multi-covering structure: given a singularity ${\cal A}_{\boldsymbol{\gamma}}$, all its integer multiples $\ell{\cal A}_{\boldsymbol{\gamma}}$, \$\0\ell\in{\mathbb{Z}}\backslash\{0\}$, appear as singularities as well. The Stokes automorphism for the singularities occurring along a half-ray $\ell{\cal A}_{\boldsymbol{\gamma}}$, $\ell\in{\mathbb{Z}}_{>0}$ is given by (3.20) (for the case in which all $c^{I}=0$) or (3.42) (for the case in which not all $c^{I}$ vanish).
@@
Figure 6: A numerical approximation to the Borel singularities of the free energies $F_{g}(t)$ in the large radius frame, for local ${\mathbb{P}}^{2}$, and for $z=-10^{-4}$. To obtain this approximation, we first consider the sequence of free energies up to $g=120$, and then we determine and plot the poles of the diagonal Padé approximant to its Borel transform. The singularity in the positive imaginary axis corresponds to (3.5) and is due to the D4 state that becomes massless at the conifold. It occurs at $2\pi{\rm i}\lambda\approx 37.5{\rm i}$. Since $2\pi t\approx 57.9+2\pi^{2}{\rm i}$, the other singularities in the plot correspond to the tower (3.11) with $d=1$, $m=0,1,-2$. As expected, singularities appear in pairs ${\cal A},-{\cal A}$.
@@ 4.1 Warm-up: the Gopakumar–Vafa duality
However, in its current form, these Gopakumar–Vafa-like dualities apply only to special toric CY geometries. We will now consider a different duality that holds conjecturally for any toric CY: the topological string/spectral theory (TS/ST) correspondence. In this correspondence, the dual of the string theory is not a field theory, but a one-dimensional quantum-mechanical model. We will end up however with matrix model representations for the topological string partition function, similar to (4.7).
@@ 4.2 Quantum mirror curves and non-perturbative partition functions
$[\mathsf{x},\mathsf{y}]={\rm i}\hbar.$ (4.8)
@@
where $\xi$ is a mass parameter of the CY, and is in principle complex. Let us point out that, in the case of mirror curves of genus $g_{\Sigma}$, there are $g_{\Sigma}$ different operators associated to $g_{\Sigma}$ different canonical forms of the curve, see cgm for a detailed explanation.
@@
Therefore, the operators $\mathsf{O}_{X}$ are functional-difference operators. Next, we will regard these operators as acting on the Hilbert space ${\cal H}=L^{2}({\mathbb{R}})$. Their domain consists of functions in ${\cal H}$ which can be extended to a strip ${\mathbb{R}}\times{\cal I}$ in the complex plane, where ${\cal I}$ is an interval depending on the operator. These functions should be square integrable along the lines ${\mathbb{R}}+{\rm i}y$, $y\in{\cal I}$.
@@ 4.3 Local ${\mathbb{P}}^{2}$, non-perturbatively
$[\mathsf{p},\mathsf{q}]=(2\pi\mathsf{i})^{-1}.$ (4.33)
@@
in terms of the variable $y$ appearing in (4.45). Implement the algorithm of bornemann to calculate $\Xi_{{\mathbb{P}}^{2}}(\kappa)$ numerically, and use this result to obtain the very first energy levels. As an example of what you should find, in Fig. 8 I plot a numerical calculation of $\Xi_{{\mathbb{P}}^{2}}(\kappa)$, obtained as follows. First, I use a Gauss–Kronrod quadrature or order $50$ to calculate the Fredholm determinant. I do this calculation for $150$ values of $\kappa$ in the interval $[-150,10]$, and I construct an interpolating function. The zeroes of this interpolating function are approximately at $-12.97004$, $-50.3105$ and $-135.882$ (rounded numerical values), in good agreement with the results in table 1 (remember the relation (4.18)). Note that the plot in Fig. 8 is indistinguishable from the plot of the same quantity that appears in mmrev . The latter was obtained by using the conjecture of ghm , which expresses the Fredholm determinant as a “quantum theta function.” The fact that they agree so precisely (within numerical errors) is an explicit test of the TS/ST correspondence. ∎
@@
(The actual planar resolvent has an additional piece, but it does not contribute to the density of eigenvalues). The density $\rho(u)$ is a symmetric one-cut distribution, and the end-points of the cut can be found from the branch cuts of the spectral curve. They are given by ±$\pm a$, where
@@
and we are assuming that $-\infty<\kappa<-3$ which corresponds to the region $-1/27<z<0$. One way of testing the result above for $\rho(u)$ is to consider a finite but large number $N$ of “typical eigenvalues” and see how they distribute along a histogram. This distribution should approximate the density of eigenvalues $\rho(u)$ as $N$ grows large. In the lectures by C. Johnson such a distribution was obtained in the case of the Gaussian matrix model by generating a random matrix with Gaussian weight, and then calculating their eigenvalues. Here the probability distribution is not Gaussian, and we will obtain the distribution in a different way, as follows. Let us consider an eigenvalue integral of the form
@@
    The discussion above makes it clear that the conjecture 4.4 can be regarded as an explicit realization of “quantum geometry,” in which the spacetime geometry of the toric CY, together with its “stringy” corrections (given by embedded Rieman surfaces) emerges from a simple quantum mechanical model on the real line. As in other string/gauge theory dualities of the ’t Hooft type, in the quantum model the CY modulus $\lambda$ is “quantized” in units of the string coupling constant. We can also think about the geometry of the CY as emerging from the eigenvalue integral (4.30) in the ’t Hooft limit. This limit is encoded in a spectral curve, which is nothing but the mirror curve we started with. Here we have seen how this works in the explicit example of local ${\mathbb{P}}^{2}$, and one can also work out the case of local ${\mathbb{F}}_{0}$ mz-wv ; zakany . We expect however this picture to hold for general toric manifolds, namely the ’t Hooft limit of the spectral traces should be described by a spectral curve given by the mirror curve. Higher order corrections to the spectral trace in the $1/N$ expansion should be governed by the topological recursion.
@@
    where the ’t Hooft parameter $t=Ng_{s}$ is identified with the flat coordinate $-{\rm i}a_{D}(u)$. This last statement can be proved rigorously bgt , providing in this way a proof of the conjecture 4.4 in this special case. The proof can be generalized to arbitrary $SU(N)$, ${\cal N}=2$ pure super Yang–Mills theories ggh . This gives further evidence for the TS/ST correspondence.
@@
    The spectral theory side guarantees that the fermionic trace $Z_{X}(N,\hbar)$ is well-defined for $N\in{\mathbb{Z}}_{>0}$, $\hbar>0$, provided the mass parameters satisfy some positivity constraints. This might be regarded as too restrictive, specially in view of e.g. approaches based on Borel resummation where it is natural to consider complex values of $\hbar$. Before filing a complaint, one should note that this is a natural aspect of dualities between strings and quantum theories, since the latter require various conditions to be well defined, and these include reality and positivity constraints on some of their parameters. However, there are indications that one can analytically continue the fermionic traces to functions with a much larger analyticity domain. One can use e.g. non–Hermitian extensions of quantum mechanics to consider complex values of $\hbar$ and of the mass parameters gm-complex ; cgum ; emr , and we expect the fermionic traces to be analytic on the cut complex plane =\$\0{\mathbb{C}}^{\prime}={\mathbb{C}}\backslash(-\infty,0]$ of the $\hbar$ variable ghm ; kama ; gm-peacock .
@@
    We have focused on relatively complicated examples, in which the mirror curve has genus $g_{\Sigma}=1$. What happens to the topological string on the resolved conifold in the context of the TS/ST correspondence, which after all is a simpler example? It turns out that the quantization of the mirror curve to the resolved conifold gives the three-term operator (4.31) with $m=1$, $n=-1$. This operator is not trace class, so strictly speaking the standard version of the correspondence does not apply in that case. Fortunately, there is a beautiful way to make sense of the spectral theory of the resolved conifold due to Y. Hatsuda hatsuda , in which one considers an appropriate analytic continuation of the spectral zeta function, instead of the fermionic spectral traces. The resulting theory leads e.g. to a spectral determinant in agreement with the general theory proposed in hmmo ; ghm .

### v3

@@ 2.1 The A model
see Fig. 1 for a pictorial representation. Let $S_{i}]\in H_{2}(M,{\mathbb{Z}})$, $i=1,\cdots,s$, be a basis for the two-homology of $M$, with $s=b_{2}(M)$ as before. The maps (2.3) are classified topologically by the homology class
@@ 2.2 The Gopakumar–Vafa representation
Note that, due to the vanishing property of the GV invariants mentioned above, the sum over $g$ in (2.19) is finite. One could think that (2.18) can perhaps be summed, as a series now in the variables ${\rm e}^{-t_{i}}$. It turns out that the properties of this series depend crucially on the value of $g_{s}$. If $g_{s}$ is real, the series is not even well-defined. This is due to the inverse square sines appearing in (2.14), which lead to singularities when $g_{s}\in 2\pi{\mathbb{Q}}$. If $g_{s}/(2\pi)$ is rational, there is a minimum degree ${\bf m}_{\rm min}$ such that infinitely many coefficients $F_{\bf m}(g_{s})$ with ${\bf m}>{\bf m}_{\rm min}$ are singular at that rational value. As a consequence, given any real value of $g_{s}$, there is a multi-degree starting from which infinitely many coefficients $F_{\bf m}(g_{s})$ can be made arbitrarily large. This is clearly a pathological situation. One could still try to make sense of the r.h.s. of (2.18) for complex values of $g_{s}$. However, in that case there is evidence that the coefficients $F_{\bf m}(g_{s})$ grow with Gargantuan speed, even worst than factorial. For example, in the case of local ${\mathbb{P}}^{2}$, introduced in (2.2), one has $b_{2}(M)=1$ (corresponding to the ${\mathbb{P}}^{1}$ inside ${\mathbb{P}}^{2}$) and therefore there is a single degree. Explicit computations suggest that hmmo ; mmrev
@@
There is a special class of toric CY manifolds which can be used to engineer five-dimensional $SU(N)$ gauge theories kkv ; n . For these manifolds, the gauge theory instanton partition function (reviewed in N. Nekrasov’s lectures in this school) provides a rearrangement of the GV expansion which is expected to converge for complex $g_{s}$ (in the case of $SU(2)$, this is proved in bs-q). However, it is still singular if $g_{s}\in 2\pi\mathbb{Q}$.
@@ 2.4 The B model
where $I,J=0,1,\cdots,s$ and $\langle\cdot,\cdot\rangle$ is the intersection pairing in $H_{3}(M^{\star})$. Integration of $\Omega$ over these cycles gives the A and B periods,
@@
The prepotential gives the genus zero free energy of the topological string, in the B model.
@@
[…]
which is a generalized Legendre transform. The function S(t,t~)S($\tilde{t}$) has the form
@@
[…]
@@
where the function S(t,t~)S($\tilde{t}$) implementing the transform is given by (2.43). The integral in the r.h.s. of (2.44) has to be understood formally, since the total free energy appearing in the integrand is itself a formal power series. To obtain the transformation properties of the genus $g$ free energies, we evaluate the integral in the r.h.s. of (2.44) in a saddle point approximation for $g_{s}$ small. At leading order we recover the generalized Legendre transform (2.39), and in particular the condition for a saddle point is precisely (2.41). The evaluation at higher orders leads to explicit transformation properties for the higher genus free energies, see abk for examples.
@@ 2.5 A more complicated example: local ${\mathbb{P}}^{2}$
Figure 2: The moduli space of local ${\mathbb{P}}^{2}$, with the three special points $z=0$ (large radius), z=1/27z=-1/27 (conifold) and $z=\infty$ (orbifold).
@@
It can be seen from the PF equations that the convergence radius of the series (2.54) at $z=0$ is set by the conifold singularity at z=1/27z=-1/27. This means in particular that the radius of convergence of the genus zero free energy (2.57) is set by the value of $t$ at the conifold point cdgp ; kz , which in this case is given by villegas
@@
The second relation is of course a consequence of (2.59). The choices of signs in (2.61) are correlated with the choice of branch cut of $\log(z)$ for 1/27<z<0-1/27<z<0, and they have been made in such a way that $\lambda$ and $\partial_{\lambda}F_{0}^{c}$ are real in that interval. The genus zero free energy in the conifold frame can be computed explicitly and it has the form,
@@ 3.1 Warm-up: the conifold and the resolved conifold
$\zeta=2\pi{\rm i}\ell\lambda,\qquad\ell\in{\mathbb{Z}}\backslash\{0\},$ (3.3)
@@
$\varphi_{\ell{\cal A}}(g_{s})={1\over 2\pi}\left\{{1\over\ell}\left({{\cal A}\over g_{s}}\right)+{1\over\ell^{2}}\right\}{\rm e}^{-\ell{\cal A}/g_{s}},\qquad\ell\in{\mathbb{Z}}\backslash\{0\}.$ (3.6)
@@
Therefore, the singularities of the Borel transform are of the form $\ell{\cal A}_{m}$, where $\{$$\ell\in{\mathbb{Z}}\backslash\{0\}$ and the “action” ${\cal A}_{m}$ is labelled by an additional integer $m$:
@@ 3.2 The general multi-instanton trans-series
[…]
@@
so (3.32) is manifestly a non-perturbative correction. The exponent in (3.35) can be interpreted as the difference between the free energies of two different backgrounds, or points in the moduli space of the CY: the background $t$, and the background $t-cg_{s}$, in which $t$ is shifted. It is easy to see that in the $\ell$-th instanton sector the shift is given by $t-\ell cg_{s}$. This is very similar to the structure of multi-instantons in Hermitian matrix models david; shenker ; msw ; multi-multi , in which the $\ell$-th instanton sector is obtained by tunnelling $\ell$ eigenvalues. Since, with the appropriate normalizations, $c$ is an integer, this suggests that $t$ has an underlying discrete structure and it is “quantized” in units of $g_{s}$. Such a quantization is typical of topological string theories with large $N$ duals, in which the CY modulus is interpreted as a ’t Hooft parameter and has the form $Ng_{s}$, where $N$ is the rank of the matrix model dv ; gv-cs ; mz . We will elaborate on this in section 4.
@@ 3.3 BPS states and the resurgent structure of topological strings
and its rank is $2(s+1)$, where we recall that $s=h^{1,2}(M)$. If we choose a basis for this lattice, we can write $\boldsymbol{\gamma}$} in terms of two pairs of vectors of rank $s+1$, with entries, ${\boldsymbol{\gamma}}=(c^{I},d_{I})$, where $I=0,1,\cdots,s$. In the context of type IIA superstring theory, the BPS states are obtained by wrapping a D2p2p brane around a cycle of even dimension $2p$ inside the CY threefold, leading to a four-dimensional BPS particle in the uncompactified directions. We can think of $c^{0}$, $d_{0}$ as D6 and D0 brane charges, respectively, and of $c^{a}$, $d_{a}$, $a=1,\cdots,s$, as D4 and D2 brane charges, respectively. The central charge corresponding to such an element of $\Gamma$ is given by
@@
where summation over the repeated indices is understood. Let us note that, in the case of toric CY manifolds, D6 branes decouple, and the charge $\boldsymbol{\gamma}$} is specified by $2s+1$ integers which we will denote by $c^{a}$, $d_{a}$ and $m$, with $a=1,\cdots,s$. The central charge reads then
@@
See Fig. 5 for a plot of the curve of marginal stability in the $u$-plane. States with charges $\boldsymbol{\gamma}$} also belong to the spectrum.
@@
    The singularities display a multi-covering structure: given a singularity ${\cal A}_{\boldsymbol{\gamma}}$, all its integer multiples $\ell{\cal A}_{\boldsymbol{\gamma}}$, $\{$$\ell\in{\mathbb{Z}}\backslash\{0\}$, appear as singularities as well. The Stokes automorphism for the singularities occurring along a half-ray $\ell{\cal A}_{\boldsymbol{\gamma}}$, $\ell\in{\mathbb{Z}}_{>0}$ is given by (3.20) (for the case in which all $c^{I}=0$) or (3.42) (for the case in which not all $c^{I}$ vanish).
@@
Figure 6: A numerical approximation to the Borel singularities of the free energies $F_{g}(t)$ in the large radius frame, for local ${\mathbb{P}}^{2}$, and for $z=-10^{-4}$. To obtain this approximation, we first consider the sequence of free energies up to $g=120$, and then we determine and plot the poles of the diagonal Padé approximant to its Borel transform. The singularity in the positive imaginary axis corresponds to (3.5) and is due to the D4 state that becomes massless at the conifold. It occurs at $2\pi{\rm i}\lambda\approx 37.5{\rm i}$. Since $2\pi t\approx 57.9+2\pi^{2}{\rm i}$, the other singularities in the plot correspond to the tower (3.11) with $d=1$, m=0,1,2m=0,1,-2. As expected, singularities appear in pairs ${\cal A},-{\cal A}$.
@@ 4.1 Warm-up: the Gopakumar–Vafa duality
However, in its current form, these Gopakumar–Vafa-like dualities apply only to special toric CY geometries. We will now consider a different duality that holds conjecturally for any toric CY: the topological string/spectral theory (TS/ST) correspondence. In this correspondence, the topological string theory is not captured by a field theory, but by a one-dimensional quantum-mechanical model. We will end up however with matrix model representations for the topological string partition function, similar to (4.7).
@@ 4.2 Quantum mirror curves and non-perturbative partition functions
$\mathsf{x},\mathsf{y}]={\rm i}\hbar.$ (4.8)
@@
where $\xi$ is a mass parameter of the CY, and is in principle complex. Let us point out that, in the case of mirror curves of genus $g_{\Sigma}$, there are $g_{\Sigma}$ different operators associated to the moduli $\kappa_{j}$, $j=1,\cdots,g_{\Sigma}$, due to the fact that there are $g_{\Sigma}$ different ways of writing the curve in the form (4.10), see cgm for a detailed explanation.
@@
[…]
@@ 4.3 Local ${\mathbb{P}}^{2}$, non-perturbatively
$\mathsf{p},\mathsf{q}]=(2\pi\mathsf{i})^{-1}.$ (4.33)
@@
in terms of the variable $y$ appearing in (4.45). Implement the algorithm of bornemann to calculate $\Xi_{{\mathbb{P}}^{2}}(\kappa)$ numerically, and use this result to obtain the very first energy levels. As an example of what you should find, in Fig. 8 I plot a numerical calculation of $\Xi_{{\mathbb{P}}^{2}}(\kappa)$, obtained as follows. First, I use a Gauss–Kronrod quadrature or order $50$ to calculate the Fredholm determinant. I do this calculation for $150$ values of $\kappa$ in the interval [150,10][-150,10], and I construct an interpolating function. The zeroes of this interpolating function are approximately at 12.97004-12.97004, 50.3105-50.3105 and 135.882-135.882 (rounded numerical values), in good agreement with the results in table 1 (remember the relation (4.18)). Note that the plot in Fig. 8 is indistinguishable from the plot of the same quantity that appears in mmrev . The latter was obtained by using the conjecture of ghm , which expresses the Fredholm determinant as a “quantum theta function.” The fact that they agree so precisely (within numerical errors) is an explicit test of the TS/ST correspondence. ∎
@@
(The actual planar resolvent has an additional piece, but it does not contribute to the density of eigenvalues). The density $\rho(u)$ is a symmetric one-cut distribution, and the end-points of the cut can be found from the branch cuts of the spectral curve. They are given by $\pm a$, where
@@
and we are assuming that $\infty<\kappa<-3$ which corresponds to the region 1/27<z<0-1/27<z<0. One way of testing the result above for $\rho(u)$ is to consider a finite but large number $N$ of “typical eigenvalues” and see how they distribute along a histogram. This distribution should approximate the density of eigenvalues $\rho(u)$ as $N$ grows large. In the lectures by C. Johnson such a distribution was obtained in the case of the Gaussian matrix model by generating a random matrix with Gaussian weight, and then calculating their eigenvalues. Here the probability distribution is not Gaussian, and we will obtain the distribution in a different way, as follows. Let us consider an eigenvalue integral of the form
@@
    The discussion above makes it clear that the conjecture 4.4 can be regarded as an explicit realization of “quantum geometry,” in which the spacetime geometry of the toric CY, together with its “stringy” corrections (given by embedded Rieman surfaces) emerges from a simple quantum mechanical model on the real line. As in other string/gauge theory dualities of the ’t Hooft type, in the quantum model the CY modulus $\lambda$ is “quantized” in units of the string coupling constant, as suggested in the discussion after (3.32). We can also think about the geometry of the CY as emerging from the eigenvalue integral (4.30) in the ’t Hooft limit. This limit is encoded in a spectral curve, which is nothing but the mirror curve we started with. Here we have seen how this works in the explicit example of local ${\mathbb{P}}^{2}$, and one can also work out the case of local ${\mathbb{F}}_{0}$ mz-wv ; zakany . We expect however this picture to hold for general toric manifolds, namely the ’t Hooft limit of the spectral traces should be described by a spectral curve given by the mirror curve. Higher order corrections to the spectral trace in the $1/N$ expansion should be governed by the topological recursion.
@@
    where the ’t Hooft parameter $t=Ng_{s}$ is identified with the flat coordinate ${\rm i}a_{D}(u)$. This last statement can be proved rigorously bgt , providing in this way a proof of the conjecture 4.4 in this special case. The proof can be generalized to arbitrary $SU(N)$, ${\cal N}=2$ pure super Yang–Mills theories ggh . This gives further evidence for the TS/ST correspondence.
@@
    The spectral theory side guarantees that the fermionic trace $Z_{X}(N,\hbar)$ is well-defined for $N\in{\mathbb{Z}}_{>0}$, $\hbar>0$, provided the mass parameters satisfy some positivity constraints. This might be regarded as too restrictive, specially in view of e.g. approaches based on Borel resummation where it is natural to consider complex values of $\hbar$. Before filing a complaint, one should note that this is a natural aspect of dualities between strings and quantum theories, since the latter require various conditions to be well defined, and these include reality and positivity constraints on some of their parameters. However, there are indications that one can analytically continue the fermionic traces to functions with a much larger analyticity domain. One can use e.g. non–Hermitian extensions of quantum mechanics to consider complex values of $\hbar$ and of the mass parameters gm-complex ; cgum ; emr , and we expect the fermionic traces to be analytic on the cut complex plane ${\mathbb{C}}^{\prime}={\mathbb{C}}\backslash(-\infty,0]$ of the $\hbar$ variable ghm ; kama ; gm-peacock . Similarly, the spectral trace is defined originally for positive integer values of $N$, leading as we mentioned before to a “discretization” of the moduli space in the quantum mechanical description. However, it is possible to extend $Z_{X}(N,\hbar)$ to an entire function of arbitrary complex values of $N$, as noted in cgm8 (at least when $\hbar>0$). Therefore, it it likely that one can obtain in the end a “smooth” description of the moduli space in terms of entire functions.
@@
    We have focused on relatively complicated examples, in which the mirror curve has genus $g_{\Sigma}=1$. What happens to the topological string on the resolved conifold in the context of the TS/ST correspondence, which after all is a simpler example? It turns out that the quantization of the mirror curve to the resolved conifold gives the three-term operator (4.31) with $m=1$, n=1n=-1. This operator is not trace class, so strictly speaking the standard version of the correspondence does not apply in that case. Fortunately, there is a beautiful way to make sense of the spectral theory of the resolved conifold due to Y. Hatsuda hatsuda , in which one considers an appropriate analytic continuation of the spectral zeta function, instead of the fermionic spectral traces. The resulting theory leads e.g. to a spectral determinant in agreement with the general theory proposed in hmmo ; ghm .